To whom may it concern,
I have trouble to get the answer for the follwoinh questions.
A bank charges 2.25% per month simple interest on personal loan. If John borrows $6400 for a period of 2 years and 1 month, find the total interest he has to pay.
Please help. Thank you very much.
CH Lam

Answer by uma(370)

Amount borrowed = $ 6400
Interest rate = 2.25%
= 0.0225
Period = 2 yrs 1 mon
= 2 1/12 yr[1 month = 1/12 yr]
= 25/12
Interest = PTR
= 6400 * 25/12 * 0.0225
= 300
Thus th3e interest = $300
Good Luck!!!